# OmniMER: Indonesian Multimodal Emotion Recognition via Auxiliary-Enhanced LLM Adaptation

### v1

@@ V-D Case Study
The textual utterance contains the Indonesian expression “galau”, which explicitly conveys negative affect and emotional distress. The text auxiliary task correctly identifies this emotion-bearing keyword and provides a strong linguistic signal for negative polarity and sadness. In contrast, the visual stream shows a slight smile and relaxed facial movements, which could be misleading if considered in isolation. The video auxiliary task captures these facial cues and interprets them as positive affect, revealing a modality-level conflict rather than suppressing it.
The audio auxiliary task analyzes prosodic patterns such as intonation and speaking rate, producing a negative-oriented interpretation. However, the acoustic evidence is relatively weak and ambiguous in this case, reflecting the inherent uncertainty of paralinguistic cues. Importantly, OmniMER does not require all auxiliary tasks to be individually correct. Instead, these modality-specific analyses collectively expose complementary and potentially conflicting emotional evidence.
@@ V-E Cross-lingual Validation
VI Conclusion
In this study, we addressed multimodal emotion recognition in low-resource language settings. We introduced IndoMER, the first multimodal emotion recognition benchmark for Indonesian, comprising 1,944 video segments from 208 speakers with aligned text, audio, and visual annotations across seven emotion categories. IndoMER presents realistic challenges including cross-modal inconsistency and long-tailed class distributions that reflect Indonesian cultural communication norms.
To tackle these challenges, we proposed OmniMER, a framework that adapts omni-modal LLMs through auxiliary modality-specific perception tasks to strengthen unimodal emotional representations prior to fusion. Experiments on IndoMER show that OmniMER outperforms the base model by 7.6 and 22.1 absolute points in Macro-F1 on sentiment classification and emotion recognition respectively, with notable gains on minority emotion categories. Cross-lingual evaluation on CH-SIMS further validates the generalizability of the framework.This work has limitations: the dataset remains relatively small, and experiments focus on a single backbone model. Future work includes extending to additional underrepresented languages, incorporating more baseline comparisons, and exploring finer-grained auxiliary objectives.

### v2

@@ V-D Case Study
The textual utterance contains the Indonesian expression “galau”, which explicitly conveys negative affect and emotional distress. The text auxiliary task correctly identifies this emotion-bearing keyword and provides a strong linguistic signal for negative polarity and sadness. In contrast, the visual stream shows a slight smile and relaxed facial movements, which could be misleading if considered in isolation. The video auxiliary task captures these facial cues and interprets them as a positive affect, revealing a modality-level conflict rather than suppressing it.
The audio auxiliary task analyzes prosodic patterns such as intonation and speaking rate, producing a negative-oriented interpretation. However, the acoustic evidence is relatively weak and ambiguous in this case, reflecting the inherent uncertainty of paralinguistic cues. Importantly, OmniMER does not require that all auxiliary tasks be individually correct. Instead, these modality-specific analyses collectively expose complementary and potentially conflicting emotional evidence.
@@ V-E Cross-lingual Validation
VI Conclusions
In this study, we addressed multimodal emotion recognition in low-resource language settings. We introduced IndoMER, the first multimodal emotion recognition benchmark for Indonesian, comprising 1,944 video segments from 208 speakers with aligned text, audio, and visual annotations across seven emotion categories. IndoMER presents realistic challenges including cross-modal inconsistency and long-tailed class distributions that reflect Indonesian cultural communication norms. To tackle these challenges, we proposed OmniMER, a framework that adapts omni-modal LLMs through auxiliary modality-specific perception tasks to strengthen unimodal emotional representations prior to fusion. Experiments on IndoMER show that OmniMER outperforms the base model in Macro-F1 on sentiment classification and emotion recognition, with notable gains on minority emotion categories. Cross-lingual evaluation on CH-SIMS further validates the generalizability of the framework.This work has limitations: the dataset remains relatively small, and experiments focus on a single backbone model. Future work includes extending to additional underrepresented languages, incorporating more baseline comparisons, and exploring finer-grained auxiliary objectives.